# AdaSCALE: Adaptive Scaling for OOD Detection

### v2

@@ 6 Conclusion
We propose AdaSCALE, a novel post-hoc OOD detection method that dynamically adjusts the scaling process based on a sample’s estimated OODness. Leveraging the observation that OOD samples exhibit larger activation shifts under minor perturbations, AdaSCALE assigns stronger scaling to likely ID samples and weaker scaling to likely OOD samples, enhancing ID-OOD separability. AdaSCALE achieves state-of-the-art performance as well as generalization across architectures requiring negligibly few ID samples, making it highly practical for real-world deployment. It lays the foundation for future research to build upon the adaptive scaling mechanism.
7 Broader Impacts
This work has positive impact in trustworthy deep learning by enabling detection of OOD samples. Reliable OOD detection can help prevent models from making overconfident predictions in high-stakes applications.
@@ D.1 Superiority of the Proposed OODness Metric ($Q^{\prime}$)
D.2 Comparison with Distance-Based OOD Detection Methods
To situate AdaSCALE within the broader landscape of OOD detection techniques, we compare its performance against representative distance-based methods, specifically NNGuide and MDS. The experiments were conducted on the large-scale ImageNet-1k benchmark using a ResNet-50 backbone. As shown in Table 13, AdaSCALE demonstrates a significant performance advantage over these methods, particularly in the challenging near-OOD setting.
@@ D.2 Comparison with Distance-Based OOD Detection Methods
D.3 Analysis on Adversarially Trained Models
@@
D.3.1 Activation Sensitivity in Adversarially Trained Networks
@@
D.3.2 Performance of AdaSCALE on Adversarially Trained Networks
@@
D.4 Robustness to Real-World Image Corruptions
@@
Appendix E Sensitivity study
E.1 Sensitivity study of $\varepsilon$
The sensitivity study of $\varepsilon$ presented at Table˜18 suggests the optimal value of $\varepsilon$ to be around 0.5.
@@ E.1 Sensitivity study of $\varepsilon$
E.2 Sensitivity study of $\lambda$
The sensitivity study of $\lambda$ presented at Table˜18 suggests the optimal value of $\lambda$ to be around 10.
@@ E.2 Sensitivity study of $\lambda$
E.3 Sensitivity study of $k_{1}$
The sensitivity study of $k_{1}$ presented at Table˜20 suggests the optimal value of $k_{1}$ to be around 1%.
@@ E.3 Sensitivity study of $k_{1}$
E.4 Sensitivity study of $k_{2}$
The sensitivity study of $k_{2}$ presented at Table˜21 suggests the optimal value of $k_{2}$ to be around 5%.
@@ E.4 Sensitivity study of $k_{2}$
E.5 Sensitivity study of $o$
We present the complete results of image perturbation study (FPR@95 $\downarrow$ / AUROC $\uparrow$) in Table˜22.
@@ E.5 Sensitivity study of $o$
E.6 Sensitivity study of $p_{\text{min}}$
We present sensitivity study (FPR@95 $\downarrow$ / AUROC $\uparrow$) of $p_{\text{min}}$, setting $p_{\text{max}}$ to 85 in the Table˜23.
@@ E.6 Sensitivity study of $p_{\text{min}}$
E.7 Joint sensitivity study
@@
The results in Table˜24 clearly indicate that the $\lambda$ hyperparameter exerts a more significant influence on the final OOD detection performance. Furthermore, a key relationship emerges: for optimal performance, a decrease in $\lambda$ must be compensated by an increase in $Q$ (and consequently, a higher $k_{1}$). This is explained by the formulation $Q^{\prime}=\lambda\cdot Q+C_{o}$, where a smaller $\lambda$ necessitates a larger $Q$ value to ensure the term $\lambda\cdot Q$ dominates the regularizer $C_{o}$.
@@
In Table˜25, when $\lambda$ is set to 0.1 with $k_{1}=1\%$, $\lambda\cdot Q$ does not dominate over $C_{o}$ term. As a result, it leads to scaling based on the regularization / balancing term instead of "predetermined" actual OOD detection signal "activation shift at peak activations". Such value of $\lambda$ (e.g., $\lambda=0.1$) invalidates the proposed hypothesis.
From the results in Table˜26, it can be observed that activation shift with all activations considered ($k_{1}=100\%$) still proves to be a useful OOD detection signal.
Furthermore, previous works (ASH, SCALE) observed that, particularly for ResNet-50 architecture, the performance drops after the percentile value of  85%. Similar observations can also be made in AdaSCALE from the results in Tables˜25 and 26.
@@
Appendix F ISH regularization:
Apart from enhancing the prior postprocessor ASH (Djurisic et al., 2023), SCALE (Xu et al., 2015) introduces a training regularization to emphasize samples with more distinct ID characteristics. We assess the performance (FPR@95 $\downarrow$ / AUROC $\uparrow$) of each method in ResNet-50 and ResNet-101 model following this regularization in Table˜27. The results indicate that AdaSCALE maintains a substantial advantage, surpassing the second-best method, SCALE, by $\textbf{12.56\%}/\textbf{5.82\%}$ and $\textbf{20.46\%}/\textbf{1.21\%}$ in FPR@95 / AUROC for near- and far-OOD detection in ResNet-50, respectively. Moreover, AdaSCALE demonstrates superior performance beyond conventional OOD detection, with corresponding improvements of 4.10%/5.87% and 9.70%/1.12% in full-spectrum setting. Furthermore, ISH regularization further amplifies the performance gap between AdaSCALE-A and OptFS, enhancing the near-OOD detection improvement from 12.96% / 6.44% to 15.18%/14.83%. These findings also generalize to ResNet-101 network.
@@ Appendix F ISH regularization:
Appendix G Hyperparameters
We present final hyperparameter values of AdaSCALE-A and AdaSCALE-L in Table˜28 and Table˜29.
@@ Appendix G Hyperparameters
Appendix H CIFAR-results
H.1 WRN-28-10
@@ H.2 DenseNet-101
Appendix I ImageNet-1k results
I.1 near-OOD detection
@@ Appendix I ImageNet-1k results
I.3 Full-Spectrum near-OOD detection
@@
I.4 Full-Spectrum far-OOD detection

### v3

@@ 6 Conclusion
We propose AdaSCALE, a novel post-hoc OOD detection method that dynamically adjusts scaling process based on a sample’s estimated OODness. Leveraging the observation that OOD samples exhibit larger activation shifts under minor perturbations, AdaSCALE assigns stronger scaling to likely ID samples and weaker scaling to likely OOD samples, enhancing ID-OOD separability. AdaSCALE achieves state-of-the-art performance as well as generalization across architectures requiring negligibly few ID samples, making it highly practical for real-world deployment. To sum up, it moves scaling-based OOD detection from heuristic-based to a more principled footing.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ D.1 Superiority of the Proposed OODness Metric ($Q^{\prime}$)
D.2 Estimated OODness $Q^{\prime}$.
Adaptive scaling depends on estimated OODness to determine the extent of scaling. We study the effect of various estimated OODness functions using ResNet-50 network (ImageNet-1k) in Table˜13. It clearly shows Q component of $Q^{\prime}$ being most critical while $\sum_{k=1}^{k_{2}}\mathbf{a}^{\varepsilon}_{\text{argsort}(\mathbf{a})_{k}}$ as correction term being a relatively superior choice. However, estimated OODness alone – without adaptive scaling – does not result in strong performance.
D.3 Comparison with Additional Baselines
@@
To situate AdaSCALE within the broader landscape of OOD detection techniques, we compare its performance against additional baselines. The experiments were conducted on the large-scale ImageNet-1k benchmark using a ResNet-50 network. As shown in Table 14, AdaSCALE demonstrates a significant performance advantage over all these methods.
ATS vs AdaSCALE: We provide a more in-depth comparison between ATS and AdaSCALE below:
  • Number of passes: AdaSCALE uses two forward passes (on the original and perturbed input) to leverage an additional independent OOD detection signal ($Q$’) in its scaling mechanism. In contrast, ATS uses only one forward pass.
  • eCDF Usage: ATS relies on multiple eCDFs, building one for each selected intermediate layer. AdaSCALE constructs a single eCDF from its final OODness score ($Q$’) in the last layer only. We hypothesize that the final layer is most effective for this, as deeper layers are responsible for semantic discrimination.
  • Role of eCDF: In ATS, eCDFs are a core part of the inference pipeline, transforming raw activations into p-values that are aggregated to compute the final scaling factor. In AdaSCALE, the single eCDF plays a more indirect, regulatory role: it calibrates the OODness score to determine an adaptive percentile, which then informs the final scaling factor calculation.
  • Data Requirement for eCDF: AdaSCALE requires only a very small number of ID samples (as few as 5-10). ATS requires access to the whole training dataset to reliably compute eCDFs for each intermediate layer.
D.4 Analysis on Adversarially Trained Models
@@
D.4.1 Activation Sensitivity in Adversarially Trained Networks
@@
D.4.2 Performance of AdaSCALE on Adversarially Trained Networks
@@
D.5 Robustness to Real-World Image Corruptions
@@
Appendix E Sensitivity study with ViT-B-16 model
Table 19: Sensitivity study of individual hyperparameters for ViT-B-16. Results are reported as FPR@95 ($\downarrow$) / AUROC ($\uparrow$).
Parameter Value Near-OOD Far-OOD
$\epsilon$
0.1 72.71 / 73.06 46.60 / 87.22
0.5 71.87 / 73.14 47.63 / 86.83
1.0 72.92 / 73.06 49.80 / 86.49
$\lambda$
1 73.92 / 72.22 47.15 / 87.15
10 71.87 / 73.14 47.63 / 86.83
20 72.94 / 73.22 51.96 / 85.93
30 73.72 / 73.06 54.55 / 84.16
$k_{1}$
1 71.87 / 73.14 47.63 / 86.83
2 72.47 / 73.37 51.60 / 85.96
3 73.20 / 73.10 54.79 / 85.12
$k_{2}$
5 71.87 / 73.14 47.63 / 86.83
10 72.06 / 73.26 47.55 / 87.03
20 72.58 / 72.87 47.89 / 86.96
$o$
2 72.28 / 73.65 47.70 / 87.02
5 71.87 / 73.14 47.63 / 86.83
10 72.34 / 73.30 48.86 / 86.69
$p_{\text{max}}$
75 76.87 / 71.78 52.97 / 86.07
80 72.60 / 73.24 48.57 / 86.78
85 71.87 / 73.14 47.63 / 86.83
90 72.40 / 73.65 48.80 / 86.64
95 72.35 / 73.63 48.73 / 86.69
Table 20: Joint sensitivity study of 4 hyperparameters in ViT-B-16. The metric reported is Validation AUROC.
$\epsilon$ $\lambda$ $o$ $p_{\text{max}}$ Val AUROC
0.1 1 2 80 87.13
85 87.55
90 87.74
5 80 87.11
85 87.53
90 87.73
10 80 87.07
85 87.47
90 87.65
10 2 80 87.46
85 87.74
90 87.78
5 80 87.43
85 87.71
90 87.77
10 80 87.41
85 87.72
90 87.79
20 2 80 87.59
85 87.77
90 87.67
5 80 87.54
85 87.73
90 87.63
10 80 87.48
85 87.68
90 87.62
0.5 1 2 80 87.08
85 87.47
90 87.65
5 80 86.94
85 87.29
90 87.41
10 80 86.65
85 86.87
90 86.88
10 2 80 87.16
85 87.47
90 87.54
5 80 86.92
85 87.25
90 87.35
10 80 86.76
85 87.15
90 87.32
20 2 80 86.86
85 86.98
90 86.86
5 80 86.40
85 86.58
90 86.53
10 80 86.09
85 86.43
90 86.55
1.0 1 2 80 87.00
85 87.34
90 87.45
5 80 86.72
85 86.90
90 86.87
10 80 85.96
85 85.81
90 85.42
10 2 80 86.73
85 87.09
90 87.22
5 80 86.31
85 86.71
90 86.93
10 80 85.97
85 86.42
90 86.69
20 2 80 86.00
85 86.23
90 86.23
5 80 85.27
85 85.60
90 85.75
10 80 84.69
85 85.09
90 85.35
Appendix F Sensitivity study with ResNet-50 model
F.1 Sensitivity study of $\varepsilon$
The sensitivity study of $\varepsilon$ presented at Table˜21 suggests the optimal value of $\varepsilon$ to be around 0.5.
@@ F.1 Sensitivity study of $\varepsilon$
F.2 Sensitivity study of $\lambda$
The sensitivity study of $\lambda$ presented at Table˜21 suggests the optimal value of $\lambda$ to be around 10.
@@ F.2 Sensitivity study of $\lambda$
F.3 Sensitivity study of $k_{1}$
The sensitivity study of $k_{1}$ presented at Table˜23 suggests the optimal value of $k_{1}$ to be around 1%.
@@ F.3 Sensitivity study of $k_{1}$
F.4 Sensitivity study of $k_{2}$
The sensitivity study of $k_{2}$ presented at Table˜24 suggests the optimal value of $k_{2}$ to be around 5%.
@@ F.4 Sensitivity study of $k_{2}$
F.5 Sensitivity study of $o$
We present the complete results of image perturbation study (FPR@95 $\downarrow$ / AUROC $\uparrow$) in Table˜25.
@@ F.5 Sensitivity study of $o$
F.6 Sensitivity study of $p_{\text{min}}$
We present sensitivity study (FPR@95 $\downarrow$ / AUROC $\uparrow$) of $p_{\text{min}}$, setting $p_{\text{max}}$ to 85 in the Table˜26.
@@ F.6 Sensitivity study of $p_{\text{min}}$
F.7 Joint sensitivity study
@@
The results in Table˜27 clearly indicate that the $\lambda$ hyperparameter exerts a more significant influence on the final OOD detection performance. Furthermore, a key relationship emerges: for optimal performance, a decrease in $\lambda$ must be compensated by an increase in $Q$ (and consequently, a higher $k_{1}$). This is explained by the formulation $Q^{\prime}=\lambda\cdot Q+C_{o}$, where a smaller $\lambda$ necessitates a larger $Q$ value to ensure the term $\lambda\cdot Q$ dominates the regularizer $C_{o}$.
@@
In Table˜28, when $\lambda$ is set to 0.1 with $k_{1}=1\%$, $\lambda\cdot Q$ does not dominate over $C_{o}$ term. As a result, it leads to scaling based on the regularization / balancing term instead of "predetermined" actual OOD detection signal "activation shift at peak activations". Such value of $\lambda$ (e.g., $\lambda=0.1$) invalidates the proposed hypothesis.
From the results in Table˜29, it can be observed that activation shift with all activations considered ($k_{1}=100\%$) still proves to be a useful OOD detection signal.
Furthermore, previous works (ASH, SCALE) observed that, particularly for ResNet-50 architecture, the performance drops after the percentile value of  85%. Similar observations can also be made in AdaSCALE from the results in Tables˜28 and 29.
@@
Appendix G ISH regularization:
Apart from enhancing the prior postprocessor ASH (Djurisic et al., 2023), SCALE (Xu et al., 2015) introduces a training regularization to emphasize samples with more distinct ID characteristics. We assess the performance (FPR@95 $\downarrow$ / AUROC $\uparrow$) of each method in ResNet-50 and ResNet-101 model following this regularization in Table˜30. The results indicate that AdaSCALE maintains a substantial advantage, surpassing the second-best method, SCALE, by $\textbf{12.56\%}/\textbf{5.82\%}$ and $\textbf{20.46\%}/\textbf{1.21\%}$ in FPR@95 / AUROC for near- and far-OOD detection in ResNet-50, respectively. Moreover, AdaSCALE demonstrates superior performance beyond conventional OOD detection, with corresponding improvements of 4.10%/5.87% and 9.70%/1.12% in full-spectrum setting. Furthermore, ISH regularization further amplifies the performance gap between AdaSCALE-A and OptFS, enhancing the near-OOD detection improvement from 12.96% / 6.44% to 15.18%/14.83%. These findings also generalize to ResNet-101 network.
@@ Appendix G ISH regularization:
Appendix H Hyperparameters
We present final hyperparameter values of AdaSCALE-A and AdaSCALE-L in Table˜31 and Table˜32.
@@ Appendix H Hyperparameters
Appendix I CIFAR-results
I.1 WRN-28-10
@@ I.2 DenseNet-101
Appendix J ImageNet-1k results
J.1 near-OOD detection
@@
J.3 Full-Spectrum near-OOD detection
@@
J.4 Full-Spectrum far-OOD detection
@@
Appendix K Visualizations of Score Separation
We present comparative visualizations of the energy score separation between SCALE and AdaSCALE using the RegNet-Y-16 architecture in terms of challenging near-OOD datasets below. As illustrated in Figure 6, AdaSCALE demonstrates a distinctively larger separation between ID and OOD score distributions compared to SCALE.
(a) ID vs ImageNet-O
(b) ID vs NINCO
(c) ID vs SSB-Hard
Figure 6: Comparative visualizations of score separation. AdaSCALE consistently shows reduced overlap between ID and OOD distributions compared to SCALE.